OpenStudy (followthefez):

If the pressure in his middle ear is normal atmospheric pressure, what is the net force exerted on his eardrum (A = 1 cm2)?

OpenStudy (followthefez):

Note, this dude is underwater. Gauge pressure at this depth underwater is P=39.2kPa Atmospheric pressure P=101.325kPa I calculated it as... P=F/A F=PA F=101.325kPa x 1cm^2 F=10.1N Note I did convert all the values to SI before calculating. This answer is incorrect. Correct answer is 3.92N. The working in the answers gives Force as being Gauge pressure multiplied by the area. Looking again at both my working, and the answer they both seem incorrect. Shouldn't the pressure be the NET pressure? ie the atmospheric pressure + the gauge pressure?

OpenStudy (anonymous):

P = F/A *plug in number 39200 Pa = F/(1 x 10 E-4 m^2) *multiply both sides by 10 E-4 F = 3.92N

OpenStudy (anonymous):

Don't forget your units when working with these kind of problems.

OpenStudy (anonymous):

Can you click best response please

OpenStudy (followthefez):

I understand the working. That's not what I was asking. I am confused about why gauge pressure is used rather than absolute pressure. Can someone please explain why this is so when there is clearly an atmospheric pressure as well as a gauge pressure acting on the eardrum?

OpenStudy (anonymous):

The problem states that "in his middle ear is normal atmospheric pressure." The problem basically asks you what the pressure would be relative to the pressure at the surface.

OpenStudy (followthefez):

So it's disregarding the fact that he's underwater?

OpenStudy (anonymous):

The net force would be the additional force he feels from being underwater.

OpenStudy (followthefez):

But the question IS asking for the net force.

OpenStudy (anonymous):

Yes, NET force. At the surface, there is 0 NET force, but under water there is.

OpenStudy (followthefez):

Exactly. So the NET force should be the sum of the two forces acting on the ear. If you only take one of the forces into account, then your not calculating the net force.